# ChronoMagic-Bench: A Benchmark for Metamorphic Evaluation of Text-to-Time-lapse Video Generation

### v1

@@ B.3 Visualization of the Different Scores of MTScore and CHScore
We provide some samples of different scoring magnitudes for MTScore and CHScore, as shown in Figure 8. It can be seen that both scores are consistent with human perception.
Appendix C More Details about ChronoMagic-Pro
C.1 Multi-Aspect Data Preprocessing
Due to the abundance of low-quality videos on video platforms, we filter out lower-quality videos based on metadata such as view counts, comments, and likes after acquiring the original videos, ultimately obtaining 66,226 original videos. Then, we use a three-stage method to further process and filter the data.
@@ C.1 Multi-Aspect Data Preprocessing
Adaptive Video Transition Cutting. In stage one: since our training data is sourced from video platforms (e.g., YouTube) where videos are designed to engage the audience, they inherently contain many transitions (significant changes in content during video playback). To address this issue, we follow the method described in Panda70M [15] to split the videos into multiple semantically consistent single-scene clips. Specifically, OpenCV [7] initially splits the video by analyzing pixel differences between adjacent frames. Let be the image frame at time ; the difference between two adjacent frames can be computed as:
@@
where and are the times of the last frame of and the first frame of , respectively. Clips are recombined where is below a certain threshold . This process results in 460K semantically consistent single-scene video clips (ChronoMagic-Pro).
Video Redundancy Elimination. In stage two: video publishers often use eye-catching titles, descriptions, or hashtags to attract traffic. As a result, time-lapse videos found through search terms may be general videos. Manually screening large-scale videos is impractical, we construct a zero-shot metamorphic-general classification strategy based on the video retrieval model [71]. Specifically, we carefully constructed 10 retrieval sentences to describe metamorphic and general videos. Let represent the -th retrieval sentence from the set of retrieval sentences . For each video in the dataset, the video retrieval model computes the relevance probability for each retrieval sentence . We denote these probabilities as . To determine if a video is general or metamorphic, we sum the relevance probabilities for each retrieval sentence. Let be the total relevance probability for general videos and be the total relevance probability for metamorphic videos:
we then judge whether the video is a time-lapse video through a voting strategy:
If , the video is classified as general; otherwise, it is classified as metamorphic. To this end, we obtain 150K video clips with higher purity and quality (ChronoMagic-ProH).
@@
Time-Aware Annotation In stage three: after obtaining high-quality time-lapse video clips, it is crucial to add appropriate captions. The simplest approach is to input the video clips into a large multimodal model to generate text descriptions of the video content. However, our experiments found that the 8B [41], 13B [76], and 34B [35] models could not accurately describe the content of time-lapse videos, resulting in severe hallucinations, as shown in Figure 10. Therefore, we decided to follow the annotation strategy of MagicTime [79]. Unlike MagicTime, due to higher costs, we adopted an open-source model [12] instead of the closed-source GPT-4V [1]. As shown in Figure 4, we first uniformly sample frames from each video segment, input these frames into the multimodal large model to describe the content, and finally have the model summarize the final video captions based on the textual descriptions of frames and the corresponding position of each frame in the video. To balance cost and effectiveness, we chose to use the 8B large multimodal model [12] instead of the 34B.
C.2 Distribution of the Generated Captions
@@
To analyze the word distribution in our generated captions within ChronoMagic-Pro, we computed their frequency distributions. The results, shown in Figure 11, reveal a prevalence of terms related to time-lapse videos, including "change," "transition," and "progressing." Additionally, words from four primary categories are evident: biological (e.g., mealworm, flower, tree), human-created (e.g., building, painting, walking), meteorological (e.g., eclipse, cloud, sunrise), and physical (e.g., burning, explosion). These terms underscore ChronoMagic-Pro’s focus on large-scale metamorphic changes, persistent transformations, and substantial physical interactions. The distribution of Captions in ChronoMagic-ProH is similar to that in ChronoMagic-Pro.
C.3 Samples of the ChronoMagic-Pro
@@
Figure 12 showcases a diverse array of samples from the ChronoMagic-Pro dataset, which features an extensive collection of time-lapse videos across several categories, including plants, buildings, ice, food, and various other objects and phenomena. Each video captures dynamic changes over time, providing rich visual information that surpasses the physical knowledge contained in many existing Text-to-Video (T2V) datasets. These samples illustrate the dataset’s diversity and depth, encompassing biological, human-created, meteorological, and physical categories, designed to support advanced research in high-dynamic text-to-video generation and related fields. Additionally, the dataset includes both time-lapse videos with significant state changes (e.g., flowers blooming) and videos with smaller state changes (e.g., clouds floating).
@@ Appendix D More Details about Experiment
D.1 Computation Resource Details
We employ two types of GPUs: NVIDIA H100 and NVIDIA A100. All implementations are conducted based on the official code using the PyTorch framework.
D.2 Details of Evaluation T2V Models
@@
Text2Video-Zero. Model Details. Text2Video-Zero [28], featuring a U-Net architecture, is a zero-shot video generation method based on the T2I model Stable Diffusion [55]. It generates latent codes for all frames using rich motion dynamics and utilizes a self-attention mechanism to enable all frames to interact with the latent codes of the first frame. This process ultimately achieves high spatial and temporal consistency in the video through denoising. It does not require training data and, therefore, does not use time-lapse videos as training data. Implementation Setups. We utilized the officially released Text2Video-Zero code and model, maintaining the original parameter settings. Specifically, we used the dreamlike-photoreal-2.0 version of Stable Diffusion [55], with a spatial resolution of 512×512 and a frame rate of 8 fps, to generate a 2-second (16-frame) video.
@@
D.3 Verification Experiment on ChronoMagic-Pro
To verify the validity and robustness of the ChronoMagic-Pro dataset, we conducted quantitative and qualitative validation experiments based on OpenSoraPlan v1.1 [40]. Specifically, we fine-tuned the temporal module of the OpenSoraPlan v1.1 model using a uniform frame extraction strategy and LoRA [17], based on the weights of OpenSoraPlan v1.1 [40]. Due to limited computational resources, we randomly selected only 10,000 video-text pairs from ChronoMagic-Pro for training. The results are shown in Table 7. After fine-tuning with ChronoMagic-Pro, the visual quality (UMT-FVD), text relevance (UMTScore), and metamorphic amplitude (MTScore and GPT4o-MTScore) were all effectively improved. Notably, the enhancement in metamorphic amplitude endowed OpenSoraPlan [40] with the ability to generate time-lapse videos of significant state changes, such as blooming flowers and city traffic. We also provide Qualitative Analysis, as shown in Figure 13. It is evident that, after fine-tuning, the generated videos can extend changes beyond mere lighting and camera movements to alterations in the state of objects, while ensuring that the visual quality, text relevance, and coherence remain uncompromised. This proves that ChronoMagic-Pro can support existing models in generating high-quality time-lapse videos with significant state changes, providing a new approach for future T2V model training. However, using a uniform frame extraction strategy and LoRA [17] for model fine-tuning may lead to decreased video coherence, which needs to be addressed in future work. In this study, we employ these methods solely for verification experiments.
@@ D.4 More Qualitative Evaluation on ChronoMagic-Bench
Due to space limitations, additional time-lapse videos generated by different baseline methods are shown in Figure 14. Similar to the results in the main text, most algorithms, except for MagicTime [79], fail to generate time-lapse videos with significant state changes, such as building construction. However, for time-lapse videos with smaller state changes, essentially faster-moving videos like city traffic changes, U-Net-based methods [66, 62, 28, 69, 22, 10, 79] exhibit much better visual quality, text relevance, and coherence compared to DiT-based methods [45, 40, 86]. This again demonstrates that U-Net-based methods are currently more stable and capable of producing satisfactory results with minimal inference. All videos generated by all models on ChronoMagic-Bench will be made publicly available.
D.5 More Quantitative Evaluation on ChronoMagic-Bench-150
We conduct a quantitative analysis of the temporal coherence of T2V models, with the results presented in Table 8. The CHScore value ranges within the same domain (open/close) are consistent, accurately reflecting the temporal coherence of different algorithms. Specifically, the temporal coherence of closed-source models is generally similar, whereas T2V-Zero [28] and OpenSora 1.2 [86] in open-source models exhibit the worst temporal coherence. The reasons are as follows: (1) Current video tracking algorithms perform poorly in detecting fluids (clouds, water flow, traffic flow), which are crucial components in time-lapse videos; (2) Closed-source models can generate realistic fluids following prompt instructions, whereas open-source models produce less realistic static fluids. Consequently, CHScore can only be adapted to current open-source models, indirectly highlighting the significant differences between open-source and closed-source models.
| Method | Venue | Backbone | Status | CHScore |
| Gen-2 [56] | Runway | U-Net | Close-Source | 5.27 |
| Pika-2.0 [33] | PikaLab | U-Net | Close-Source | 4.00 |
| Dream Machine [44] | LUMA | DiT | Close-Source | 2.30 |
| KeLing [32] | Kwai | DiT | Close-Source | 3.69 |
| ModelScopeT2V [66] | Arxiv’23 | U-Net | Open-Source | 10.64 |
| ZeroScope [62] | CVPR’23 | U-Net | Open-Source | 24.10 |
| T2V-zero [28] | ICCV’23 | U-Net | Open-Source | 1.84 |
| LaVie [69] | Arxiv’23 | U-Net | Open-Source | 9.58 |
| AnimateDiff [22] | ICLR’24 | U-Net | Open-Source | 11.09 |
| VideoCrafter2 [10] | Arxiv’24 | U-Net | Open-Source | 7.78 |
| MCM [80] | Arxiv’24 | U-Net | Open-Source | 14.14 |
| MagicTime [79] | Arxiv’24 | U-Net | Open-Source | 11.58 |
| Latte [45] | Arxiv’24 | DiT | Open-Source | 13.79 |
| OpenSora 1.1 [86] | Github’24 | DiT | Open-Source | 10.46 |
| OpenSora 1.2 [86] | Github’24 | DiT | Open-Source | 5.60 |
| OpenSoraPlan v1.1 [40] | Github’24 | DiT | Open-Source | 10.32 |
D.6 Details of Human Evaluation
@@
To validate the effectiveness of the automated metrics, we selected a subset of videos for user evaluation, inviting 171 participants to provide manual evaluation results. To enhance user satisfaction, we chose five representative baseline results (AnimateDiff [22], MagicTime [79], VideoCrafter2 [10], Opensora 1.1 [86], OpenSoraPlan v1.1 [40]) for users to choose from. Following established methodologies from prior studies [53, 65, 79, 59], we designed a detailed questionnaire for human evaluators to rate the generated content. The evaluation focused on four primary aspects: Visual Quality, Text Relevance, Metamorphic Amplitude, and Coherence. For each criterion, we employed a five-point rating scale and provided scoring guidelines to ensure consistent user selections, thereby minimizing assessment bias. For detailed criteria, please refer to Figure 15.
@@ Appendix E More Details about 75 Subcategories in ChronoMaigc-Bench
Station. Focuses on the activities within transportation stations, such as train stations, bus terminals, and airports. These videos capture the flow of passengers, arrivals, departures, and the hustle and bustle of travel hubs.
Traffic. Captures the movement of vehicles on roads and highways, including the traffic flow, congestion, and the changing pace of vehicular movement throughout day.
@@
Beach. Illustrate the changes in tides, waves, and shifting weather conditions throughout day.
@@
Appendix F Licensing, Hosting and Maintenance Plan
Author Statement. We take full responsibility for the licensing, distribution, and maintenance of our ChronoMaigc-Bench and ChronoMagic-Pro.
License. ChronoMaigc-Bench and ChronoMagic-Pro are under CC-BY 4.0 license.
Hosting. The code and dataset are uploaded to GitHub and HuggingFace and made public. The dataset is in the JSON file format.
Metadata. Metadata and Benchmark are uploaded to Huggingface https://huggingface.co/spaces/BestWishYsh/ChronoMagic-Bench.

### v2

@@ B.4 Visual Reference of the Different Scores of MTScore and CHScore
We also provide some samples of different scoring magnitudes for MTScore and CHScore, as shown in Figure 8. It can be seen that both scores are consistent with human perception. We strongly recommend checking out the Project Page, which provides more case studies on the metrics.
Appendix C More details about ChronoMaigc-Pro
C.1 Data Preprocessing
@@
Due to the abundance of low-quality videos on video platforms, we filter out lower-quality videos based on metadata such as view counts, comments, and likes after acquiring the original videos, ultimately obtaining 66,226 original videos. Additionally, since our training data is sourced from video platforms (e.g., YouTube) where videos are designed to engage the audience, they inherently contain many transitions (significant changes in content during video playback). To address this issue, we follow the method described in Panda70M [16] to split the videos into multiple semantically consistent single-scene clips. Specifically, OpenCV [8] initially splits the video by analyzing pixel differences between adjacent frames. Let be the image frame at time ; the difference between two adjacent frames can be computed as:
@@
where and are the times of the last frame of and the first frame of , respectively. Clips are recombined where is below a certain threshold . This process results in semantically consistent single-scene video clips.
C.2 Time-Aware Annotation
@@
After obtaining high-quality time-lapse video clips, it is crucial to add appropriate captions. The simplest approach is to input the video clips into a large multimodal model to generate text descriptions of the video content. However, our experiments found that the 8B [42], 13B [79], and 34B [36] models could not accurately describe the content of time-lapse videos, resulting in severe hallucinations, as shown in Figure 10. Therefore, we decided to follow the annotation strategy of MagicTime [83]. Unlike MagicTime, due to higher costs, we adopted an open-source model [13] instead of the closed-source GPT-4V [1]. As shown in Figure 10, we first uniformly sample frames from each video segment, input these frames into the multimodal large model to describe the content, and finally have the model summarize the final video captions based on the textual descriptions of frames and the corresponding position of each frame in the video. To balance cost and effectiveness, we chose to use the 8B multimodal large model [13] instead of the 34B.
C.3 Distribution of the Generated Captions
@@
To analyze the word distribution in our generated captions within ChronoMagic-Pro, we computed their frequency distributions. The results, shown in Figure 11, reveal a prevalence of terms related to time-lapse videos, including "change," "transition," and "progressing." Additionally, words from four primary categories are evident: biological (e.g., mealworm, flower, tree), human creation (e.g., building, painting, walking), meteorological (e.g., eclipse, cloud, sunrise), and physical (e.g., burning, explosion). These terms underscore ChronoMagic-Pro’s focus on large-scale metamorphic changes, persistent transformations, and substantial physical interactions.
C.4 Samples of the ChronoMagic-Pro
@@
Figure 12 showcases a diverse array of samples from the ChronoMagic-Pro dataset, which features an extensive collection of time-lapse videos across several categories, including plants, buildings, ice, food, and various other objects and phenomena. Each video captures dynamic changes over time, providing rich visual information that surpasses the physical knowledge contained in many existing Text-to-Video (T2V) datasets. These samples illustrate the dataset’s diversity and depth, encompassing biological, human creation, meteorological, and physical categories, designed to support advanced research in high-dynamic text-to-video generation and related fields. Additionally, the dataset includes both time-lapse videos with significant state changes (e.g., flowers blooming) and videos with smaller state changes (e.g., clouds floating).
C.5 Additional Statements
1. The aesthetic detector exhibits inherent biases, favoring artistic images, such as oil paintings and other art forms, over more realistic styles. Consequently, low aesthetic scores do not necessarily indicate poor-quality data. Retaining a small portion of such data can enhance the diversity of the videos. Thus, we include 13% of clips with low aesthetic scores in ChronoMagic-Pro.
2. There are two types of time-lapse videos: compressed and uncompressed. The former represents the entire process in a few seconds, while the latter can last for several minutes or even tens of minutes. ChronoMagic consists of compressed videos, whereas ChronoMagic-Pro includes both types to increase diversity. If the 60s+ videos, which account for 27% of ChronoMagic-Pro, are excluded, the average length would be only 12.36 seconds.
3. Since the data of ChronoMagic-Bench and ChronoMagic-Pro both include videos from YouTube, we deduplicate the data using video IDs. Additionally, we employed different annotation models (e.g., GPT-4o [1], ShareGPT4Video [13]) to label the benchmark and dataset, further reducing the risk of data leakage.
@@ Appendix D More Details about Experiment
D.1 Details of Resource
We employ two types of GPUs: Nvidia H100 (x8) and Nvidia A800 (x8). All implementations are conducted based on the official code using the PyTorch framework.
D.2 Details of Evaluation Models
Since most T2V models do not support dynamic resolution or variable duration, it is not feasible to standardize these parameters. Therefore, we follow the official popular settings [45, 26, 61, 75] to maintain a degree of fairness. Moreover, both MTScore and CHScore mitigate the influence of resolution by employing a resizing strategy that adjusts the shorter edge and utilizes center cropping. MTScore further employs a fixed frame extraction method to ensure a consistent frame count, while the different terms of CHScore are insensitive to num_frames, thereby mitigating discrepancies due to varying frame numbers.
@@ D.2 Details of Evaluation Models
T2V-Zero. Model Details. Text2Video-Zero [28], featuring a U-Net architecture, is a zero-shot video generation method based on the T2I model Stable Diffusion [57]. It generates latent codes for all frames using rich motion dynamics and utilizes a self-attention mechanism to enable all frames to interact with the latent codes of the first frame. This process ultimately achieves high spatial and temporal consistency in the video through denoising. It does not require training data and, therefore, does not use time-lapse videos as training data. Implementation Setups. We utilized the officially released Text2Video-Zero code and model, maintaining the original parameter settings. Specifically, we used the dreamlike-photoreal-2.0 version of Stable Diffusion [57], with a spatial resolution of 512×512 and a frame rate of 8 fps, to generate a 2-second (16-frame) video.
@@
CogVideoX Model Details. CogVideoX [81] is a state-of-the-art text-to-video diffusion model that builds upon the success of large-scale DiT models. To enhance text-video alignment, CogVideoX utilizes an expert transformer with expert adaptive LayerNorm, facilitating deep fusion between modalities. The model implements 3D full attention to comprehensively model videos along both temporal and spatial dimensions, ensuring temporal consistency and capturing large-scale motions. Its training data consists of scraped videos and images, and custom refined Panda70M [16], COCO caption [43] and WebVid [2]. Implementation Setup. We use the officially released CogVideoX code and model. For our experiments, we set the spatial resolution to 720x480, generated 48 frames, and used a frame rate of 8 fps, resulting in a 6-second video.
EasyAnimate Model Details. EasyAnimate [77] is an advanced text-to-video generation model designed to create high-quality animated videos from textual prompts. It adopts U-ViT [3] architectures and slice-vae to avoid unstable training. Its training data consists of videos and images scraped from open-source websites, and open-source dataset 10M SAM [30] and 2M JourneyDB [65]. The official documentation does not specify that time-lapse videos were used. Implementation Setup. We utilized the officially released EasyAnimateV3 code and model. For our experiments, we used the 720P version of the model. As per the default setting, we set the spatial resolution to 1008x576, generated 96 frames, and used a frame rate of 24 fps, resulting in a 4-second video.
D.3 Further Verification Experiment on ChronoMagic-Pro
Notably, after fine-tuning in ChronoMagic-Pro, the enhancement in metamorphic amplitude endowed OpenSoraPlan [41] with the ability to generate time-lapse videos of significant state changes, such as blooming flowers and city traffic. We provide additional qualitative analysis, as shown in Figure 13. It is evident that, after fine-tuning, the generated videos can extend changes beyond mere lighting and camera movements to alterations in the state of objects, while ensuring that the visual quality, text relevance, and coherence remain uncompromised. This proves that ChronoMagic-Pro can support existing models in generating high-quality time-lapse videos with significant state changes, providing a new approach for future T2V model training. Moreover, our findings suggest that with appropriate fine-tuning, it is possible to correct the common tendency of video models to produce nearly static videos on arbitrary topics. This phenomenon has also been observed in MagicTime-DiT [83], despite utilizing only around 2,000 time-lapse videos. However, it is important to note that the Magic Training Strategy [83], originally designed for U-Net-based models, may not be as effective for DiT-based models. In this study, we employ this methods solely for verification experiments. Additionally, the efficacy of simple fine-tuning is inferior to that achieved through the Magic Training Strategy [83].
@@ D.4 More Qualitative Evaluation on ChronoMagic-Bench
Due to space limitations, additional time-lapse videos generated by different baseline methods are shown in Figure 14. Similar to the results in the main text, most algorithms, except for MagicTime [83], fail to generate time-lapse videos with significant state changes, such as building construction. However, for time-lapse videos with smaller state changes, essentially faster-moving videos like city traffic changes, U-Net-based methods [68, 64, 28, 71, 22, 11, 83] exhibit much better visual quality, text relevance, and coherence compared to DiT-based methods [47, 41, 90]. This again demonstrates that U-Net-based methods are currently more stable and capable of producing satisfactory results with minimal inference. All videos generated by all models on ChronoMagic-Bench is publicly available on https://pku-yuangroup.github.io/ChronoMagic-Bench.
D.5 More Analysis of Closed-Source Models
We present and analyze the results from a qualitative perspective, as shown in Figure 15. The results are consistent with Table 4. For metamorphic amplitude, most methods can only generate simple time-lapse videos, such as traffic flow; only Dream Machine [46] can generate a moderately challenging full process of night-to-day transformation; no method can generate complex changes like plant growth or building construction. In terms of temporal coherence, the performance of various closed-source models is comparable, with minor visible differences. Regarding visual quality, the DiT-based methods Dream Machine [46] and KeLing [33] outperform those based on U-Net, producing more realistic plants, more accurately saturated sky colors, and clearer traffic flow. In terms of text relevance, all methods adhere to the prompt’s instructions to generate content relevant to the theme, except for Pika-2.0 [34], which mistakenly interprets day-to-night as night-to-day.
D.6 Additional Details of Human Evaluation
@@
Pre-processing The questionnaire for human evaluators to rate the generated content was established following methodologies from prior studies [55, 67, 83, 61]. The evaluation focused on four primary aspects: Visual Quality, Text Relevance, Metamorphic Amplitude, and Coherence. For each criterion, we employed a five-point rating scale and provided scoring guidelines to ensure consistent user selections, thereby minimizing assessment bias. For detailed criteria, please refer to Figure 17. For detailed explanation of voters, the voter population predominantly comprises undergraduate, master’s, and phd students from universities, along with a segment of the general public who are not associated with this field. They come from various regions around the world, including China, USA, Singapore, etc., which ensures that the participants have universality. This composition guarantees the precision and diversity of human evaluations.
Pose-processing Given the use of a simple five-point evaluation scale, we remove outliers from the responses as follows:
Restricted each IP address to prevent duplicates and required users to log in to their accounts before voting, ensuring that each person could only submit once.
Determined the validity of data based on the time taken to complete the questionnaire. Given that completing a questionnaire typically takes 10 to 20 minutes, we excluded samples where the response time was less than 10 minutes.
Randomized the order in which different videos were presented to avoid cognitive biases among voters.
Required a sliding verification at submission to ensure that all questionnaires were completed manually and not by bots.
Discarded any questionnaire where 50% of the ratings were extreme values, i.e., the sum of 5-point and 1-point options exceeded 50%.
Additional Evaluation In addition to the main text, Figure 6 analyzes the video metrics aggregated by the model. We also provide a human evaluation of disaggregated data (that is, where each point represents a video), which consists of 32 videos randomly selected from all the questionnaires. The results are shown in Figure 16. It can be seen that the proposed MTScore and CHScore are consistent with human perception in terms of disaggregated data.
@@ Appendix E More details about 75 subcategories in ChronoMaigc-Bench
Station. Focuses on the activities within transportation stations, such as train stations, bus terminals, and airports. These videos capture the flow of passengers, arrivals, departures, and the hustle and bustle of travel hubs..
Traffic. Captures the movement of vehicles on roads and highways, including the traffic flow, congestion, and the changing pace of vehicular movement throughout the day.
@@
Beach. Illustrate the changes in tides, waves, and shifting weather conditions throughout the day.
@@
Appendix F Ethics Statement
Potential Harms Caused by the Research Process. The video data utilized by ChronoMagic-Bench is sourced from free content available on four platforms: Pexels (CC0), MixKit (CC0), PixaBay (CC0), and YouTube (CC BY 4.0). Conversely, ChronoMagic-Pro exclusively employs videos from YouTube (CC BY 4.0). The licensing types of these videos are clearly indicated on their respective platforms. The CC0 license (Creative Commons Zero) designates content as public domain, allowing unrestricted use without the need for additional permissions or licenses. Videos from the YouTube platform adhere to the CC BY 4.0 license (Creative Commons Attribution 4.0); consequently, we have included video IDs and author information in the metadata to prevent any potential contractual disputes. The video content consists entirely of time-lapse footage, and we detect and discard NSFW content based on the video caption. For videos involving identifiable individuals, we accelerate the blurring process to ensure the security of personally identifiable information. The collected videos are organized into four major categories (comprising 75 subcategories), with contributors hailing from various countries and regions worldwide. This diversity ensures that ChronoMagic-Bench and ChronoMagic-Pro possess ample representativeness. The Open-Sora-Plan model [41], fine-tuned using our dataset, exhibited no significant content bias.
Data collection was facilitated by the dedicated efforts of numerous contributors, including the authors of this paper and those who participated in the human evaluation. We regard an individual’s hourly wage or compensation as personal information, which, due to privacy considerations, cannot be disclosed. Nonetheless, we can confirm that all participants received appropriate compensation in compliance with the legal requirements of their respective countries or regions. The privacy information of all participants is protected, so there is no additional risk to the them.
Societal Impact and Potential Harmful Consequences. The objective of ChronoMagic-Bench is to identify the limitations of current text-to-video generation models in producing time-lapse videos and to develop the ChronoMagic-Pro dataset to advance the field. Although time-lapse video generation models offer substantial potential to support and enhance human creativity, it is crucial to consider broader societal implications during their development:
First and foremost, environmental issues cannot be overlooked. As text-to-video generation technology advances, the demand for computational resources escalates. Large-scale data processing, model testing, and training generally depend on energy-intensive data centers, which significantly contribute to carbon emissions. For instance, this study utilized the energy-intensive H100 for experiments. If not addressed, the widespread adoption of this technology could further exacerbate climate change. Consequently, researchers and developers should focus on optimizing algorithms to reduce energy consumption.
Secondly, the generation of false content has raised significant social and ethical concerns. After appropriate fine-tuning using the ChronoMagic-Pro/ProH dataset, text-to-video generation models are capable of producing not only metamorphic videos with extended time spans and high levels of realism but also high-quality general videos. These generative models could be misused to create deceptive videos, potentially misleading the public or disseminating misinformation, particularly on fast-paced and widely influential platforms such as social media. To prevent such misuse, it is crucial to consider the implementation of content authenticity verification mechanisms and the establishment of robust legal and ethical frameworks during the development and deployment of these technologies.
Lastly, the issue of dataset bias may result in skewed and inequitable outcomes in model generation. Although the video content in the ChronoMagic-Bench and ChronoMagic-Pro datasets is sourced globally, the captions are exclusively in English. This single-language choice may impair the model’s ability to accurately interpret and generate videos across diverse cultural contexts and non-English language environments. Moreover, this bias could exacerbate existing language inequalities by disregarding the needs of non-English-speaking users. Therefore, future dataset construction should incorporate multilingual support to ensure broader adaptability and fairness in models on a global scale.
Impact Mitigation Measures. We take full responsibility for the licensing, distribution, and maintenance of our ChronoMaigc-Bench and ChronoMagic-Pro/ProH. Our datasets and benchmark are released under a CC-BY-4.0 license, and our code under an Apache license. We have clearly stated on our homepage that all data is for academic research only to prevent misuse or improper use. And we provide the email address for YouTube authors to contact and remove invalid videos in time. All the metadata are hosted on GitHub and HuggingFace at the following URLs: https://github.com/PKU-YuanGroup/ChronoMagic-Bench and https://huggingface.co/collections/BestWishYsh.